Thank you in advance for helping me with the following problems. I am struggling with these and they are due tonight by midnight. So any help given would be appreciated.
We are working with quadratic equations, functions and inequalities with graphing.
1. 5x - 10x^2 < 0 Solve each inequality, state the solution set using interval notation and graph the solution set.

Answer by stanbon(75887)

Solve the equality 1.5x-10x^2=0
Multiply by 10 to get"
15x- 100x^2=0
Factor to get x(15 - 100x)=0
x=0 or x= 0.15
Plot these two value on a number line thus forming three
intervals: x<0, 00.15
Check a value in each of these intervals to see if all
values in the interval are and or not part of the solution
to the inequality.
For example:
If =-10 you get 1.5(-10)-10(-10)^2<0 ?
-15-1000<0 This is true so all x values
less than 0 are part of the solution.
Do the same for x = 0.15/2 and x=10
The same procedure should be followed on your
other posted problems.
Cheers,
Stan H>
Cheers,
S